# A Survey of the Self Supervised Learning Mechanisms for Vision Transformers

### v5

@@ 3.1.4 Hybrid Methods
Hybrid methods similar to generative approaches often combine the strengths of different techniques to achieve specific objectives [104].
@@
Tianhu et al. (2025) [105], proposed a hybrid deep learning model MAEMC-NET, which combines masked autoEncoders for reconstructing masked CT patches and Momentum Contrast (MoCo) for learning discriminative features from different CT views.
Wu et al. (2024) [106] implemented the hybrid SSL pre-training framework Enhanced SSL with Masked Autoencoders (ESLA). This innovative approach integrates both contrastive learning (CL) and MAE frameworks to enhance SSL. ESLA improves model generalization and representation while addressing issues stemming from competing features during data enhancement. This is achieved through a masking mechanism in the CL framework and by blending representations of different views in the latent space.
@@
Quan et al. (2023) [107] introduced a novel SSL model, based on the global contrast-masked autoencoder (GCMAE), which adeptly extracts features from pathological images including both local and global features. GCMAE leverages masking image reconstruction and contrastive learning as self-supervised pretext tasks, enabling the encoder part to proficiently depict local and global features. This model is comprises of an initializer, encoder, a tile feature extractor, and a global feature extractor facilitating image reconstruction and contrastive learning tasks. In this investigation, an asymmetric encoder-decoder architecture is employed. ViT serves as the backbone for the encoder, pretrained using MAE. The module responsible for extracting features from patches consists of 8 transformer blocks, creating an asymmetric configuration when combined with an encoder comprising approximately 12 transformer blocks, as seen in the ViT-base architecture. This asymmetry facilitates decoupling of the encoder and decoder, fostering the acquisition of more generalized representations. Extraction of global feature is achieved via utilizing contrastive learning. The hidden representation Vpatch(vis) is used to reconstruct image through decoder and update a memory bank (B). This updation of B is executed to store the global features utilizing momentum coefficient t. The dynamically update is conducted to pursue the feature vectors of image data using memory bank as a fixed size queue. The features randomly selected from memory bank are utilized as negative samples, that helps to reduce the effects of the constraint caused by batch size on the heavy performance Notably, the momentum adjustment mechanism is designed distinctively from MoCo’s model parameter update, it employ an encoder structure and a memory bank to facilitate contrastive learning. Vpath(vis) is depicted as the encoders output in the current epoch in the latent representation, where as the vpatch(vis) represented the previous epoch latent representation stored in B. Particularly, the latent demonstrations at the output of encoder as vpatch(vis) in the current epoch, The cost function combines both, the weighted sum of MSE loss for tile feature extraction and the NCE loss for global feature extraction. This formulation aims to minimize the gap between similar features, enhancing the model’s generalization and accuracy, particularly in cross-dataset transfer learning tasks.
@@
Fang et al. (2023) [108] proposed a hybrid SSL framework (HSL) for Hyperspectral Image (HSI) classification. Leveraging MIM and contrastive learning to enhance its effectiveness. Employing a ViT as the backbone network within an asymmetric encoder-decoder architecture with two branches, HSL achieve improved HSI classification utilizing contrastive pre-training and self-supervised generative tasks by effectively extracting spatial and spectral details. To facilitate image reconstruction through masked inputs, each branch connects intermediate features of the encoder to the decoder at the corresponding level. Contrastive learning is accomplished by utilizing the intermediate features that are obtained from two encoders that share parameters across both branches. Utilizing standard data augmentation methods, HSL generates several views of input images. Then each image view undergoes random masking technique such as MAE. Initially, image segments with mask are reorganised into a series of patches and later processed by HSL encoder and decoder to extract patterns. Extracted features are then pass through an adaptive average pooling layer and a projection head to achieve the contrastive learning task of capturing image resemblance. The decoder, which also includes multiple Transformer encoders, rebuilds patches after certain patches have been masked. Moreover, the Transformer Encoder retains the initial structure while the two output features from the encoder are used to compute contrastive loss. The loss is calculated using the decoded sequence produced by the decoder and the sequence that precedes the encoder masking, efficiently completing the masked image rebulding process. In summary, HSL combines contrastive pre-training and self-supervised generative tasks, facilitating effectiveness for downstream HSI classification task.
@@ 4.1 BYOL
Bootstrap Your Own Latent (BYOL) [38] employs a teacher-student framework, where the student network predicts the output of a teacher network on augmented inputs. The teacher network is updated as a moving average of the student’s parameters, ensuring learning consistency. This method mirrors the stabilization effect seen in regularization techniques like batch normalization and parameter sharing.
@@ 4.2 MoCo
Momentum Contrast (MoCo) [31] leverages a momentum-based moving average of the query encoder, establishing a large and consistent dictionary. This approach acts as a dynamic regularization mechanism, ensuring stability in the representation of dictionary keys over time. This stability is vital for the extraction of stable and generalizable features, akin to how traditional regularization methods prevent overfitting by maintaining consistency in model parameters.
@@ 4.3 SimCLR
SimCLR [32] utilizes a straightforward contrastive learning framework, where its regularization is primarily through aggressive data augmentation techniques. These augmentations consist of random cropping, resizing, color manipulation, and the application of Gaussian blur, enabling the model to learn invariant and robust features. This is similar to how techniques like dropout and data augmentation in traditional settings prevent over-reliance on specific input features.
@@ 4.4 DINO
SINO (Self-Distillation with No Labels) [61] utilizes self-distillation for regularization, where the student network mimics the teacher network, updated as an exponential moving average of the student’s parameters. This method is reminiscent of techniques like early stopping and weight decay, which aim to prevent overfitting by smoothing the learning trajectory.
@@ 4.5 Barlow Twins
Barlow Twins [65] method aims to align the cross-correlation matrix between the outputs of twin networks with the identity matrix. This alignment encourages the model to learn features invariant to input distortions, functioning as a regularization technique by discouraging the learning of redundant features.
@@ 4.6 SwAV
SwAV (Swapping Assignments between Views) [37] regularization is unique, involving clustering approaches and consistency enforcement between cluster assignments for alternate projections of the single image. This approach is akin to enforcing a constraint (regularization) on the feature space to enhance generalizability.
@@ 4.7 TiCo
TiCo (Time-Contrastive Consistent Learning of Visual Representations) [78] focuses on learning temporally consistent representations, acting as a regularization mechanism by ensuring robustness to temporal variations in data. This is similar to entropy regularization, which encourages the model to learn more uncertain, hence generalizable, features.
@@ 4.8 CMC
CMC (Contrastive Multiview Coding) [79] regularizes learning by enforcing invariance to different views of the same scene. This invariance is a form of regularization, similar to dropout variations that encourage the model to learn features that are invariant to specific changes in the input.
@@ 4.9 SELF-LABELLING
SELF-LABELLING [76] approach involves pseudo-labeling, where the network assigns labels to its inputs and learns from these self-generated labels. This process regularizes the model by forcing it to generalize from its initial, potentially noisy, self-generated labels, akin to label smoothing which also deals with label noise.
@@ 5.2.1 ImageNet
The ImageNet classification dataset [109] contains roughly 1.2 million images for training, fifty thousand images for validation, and 100000 images for testing. Overall, the dataset is divided into one thousand diverse object classes including tablelamp, goldfish, radiator, screwdriver, and so on.
@@ 5.2.2 COCO
COCO (Common Objects in Context) [110] is a large-scale object detection, key-point detection, segmentation, and captioning dataset. The dataset contains over 330,000 images of which around 200,000 are fully labeled. The dataset is divided into 80 object classes and over 90 generic ‘stuff’ classes.
@@ 5.2.3 CIFAR-10
The CIFAR-10 a short for Canadian Institute For Advanced Research is a dataset [111] that contains sixty thousand images divided into 10 classes in which every class contains 6000 images. In total, 50000 images are utilized for training purpose while the rest of 10000 images are utilized for testing purpose. The dataset comprises of 10 classes, including horse, airplane, frog, automobile, cat, deer, bird, dog, ship, and truck.
@@ 5.2.4 CIFAR-100
CIFAR-100 dataset [111] is an extension of CIFAR-10. Although the total number of images are same on both datasetss, i.e., sixty thousand, CIFAR-100 is divided into 100 distinct classes with 600 images per class. Similar to CIFAR-10, a total of 50000 images are utilized for training and 10000 images are employed for testing. Every image contains two labels, a coarse label (general category) and a fine label (particular class). Some particular object classes include apple, bee, cattle, chair, rabbit, and wowan, whereas general categories include flowers, household_furniture, medium_mammals, people, etc.
@@ 5.3.1 InfoNCE Loss
InfoNCE loss (Information Noise Contrastive Estimation) [60] is a loss function commonly used in contrastive learning, particularly in self-supervised learning tasks. It operates by introducing noise into the input data to create negative pairs, enabling the model to differentiate between real positive pairs and artificially generated negatives.
@@ 5.3.2 NT-Xent/SimCLR Loss
NT-Xent (Normalized Temperature Scaled Cross Entropy) [32], which extends InfoNCE by incorporating a temperature parameter to scale the pairwise similarities. Through a softmax function, it computes the negative log-likelihood of positive pairs, enhancing the overall contrastive learning process.
@@ 5.3.3 Triplet Loss
In many machine learning tasks, particularly in face recognition, person identification, and similarity-based retrieval, a key challenge is learning meaningful feature representations that ensure similar instances are clustered together while dissimilar ones are well separated. Traditional classification-based losses struggle in scenarios with high intraclass variability and low interclass differences, leading to misclassifications. Triplet loss [112] addresses this issue by using a relative distance-based approach, which ensures that an anchor sample is closer to a positive sample (the same class) than to a negative sample (different class) by a specified margin. By simultaneously minimizing the anchor-positive distance and maximizing the anchor-negative distance, Triplet Loss enforces a structured embedding space that enhances discrimination. This approach has proven highly effective in deep metric learning tasks, such as FaceNet, where learning compact and well-separated embeddings is crucial for robust face verification, clustering, and retrieval in unseen data scenarios.
@@ 5.3.6 Circle Loss
In response to the shortcomings of triplet loss, the Circle Loss[113] was recently developed. By establishing a decision boundary in a hypersphere around instances and incorporating a flexible buffer based on the instance distances, it encourages positive pairs to lie within the boundary while keeping negative pairs outside.
These are just a few instances of the wide array of contrastive loss functions employed in various research works. The selection of a loss function is determined by factors like the specific problem, the nature of the data, and the required characteristics of the learned representations. Researchers often experiment with multiple loss functions to determine the most suitable one for their specific task [114]. As contrastive learning progresses, more innovative loss formulations and improvements are likely to emerge, opening up new possibilities for unsupervised representation learning .
@@ 5.4 Comparison with Supervised Learning Approaches
SSL, particularly in the context of ViTs, has completely changed the area of computer vision. Large volumes of unlabeled data can be used with these techniques to acquire meaningful representations, reducing the substantial reliance on labeled instances observed in conventional supervised learning. The main benefit of SSL is its capacity to take advantage of the wealth of unlabeled data that is readily available online [32], allowing models to learn from big datasets without the need for explicit annotations. SSL is especially appealing for real-world applications because of its scalability and affordability given that manual data labeling can be resource-intensive, costly, and require significant effort [6].
The comparative evaluation of SSL and supervised learning methods using the same downstream tasks and evaluation metrics has shed light on the effectiveness of SSL in ViTs. In situations where the annotated data is limited or challenging to gather, SSL has demonstrated its potential to achieve competitive or even superior performance. This is especially valuable for real-world scenarios where manually annotating data might not be feasible, yet high-performance computer vision models are needed. A comprehensive learning framework known as ”SSL” employs pretext tasks derived solely from unsupervised data. These pretext tasks are designed to learn the meaningful visual representation in order to complete them effectively [32]. Whereas, in the supervised learning paradigm, each input sample is connected to its matching target or output label, and the system is trained on a labeled dataset. In order for the model to accurately predict outcomes on data that has not yet been observed, supervised learning aims to establish a relationship between input data and their associated output labels [115].
@@ 6 Comparative Analysis of SSL Mechanisms
In this section, we will provide a comparative analysis of state-of-the-art SSL (SSL) mechanisms employing ViTs. Various SSL methods used different networks as their backbone networks. For an unbiased comparison, we compare different SSL methods employing ViT-B/16 [11] as their backbone network.
@@ 6.1 Performance Comparison
We compare SOTA SSL methods, in Table 4, where self-supervised pre-training is performed over ImageNet-1K [109] training dataset. Later, fully supervised training is performed to evaluate the feature representations using (i) linear probing and (ii) end-to-end fine-tuning. The inputs are resized to 224x224 crops and report top-1 validation accuracy. From Table 4, we notice that DINO [38], MoCo-V3 [93], MSG-MAE [116], iBOT [117], and EsViT [118] achieve more than 75% top-1 accuracy, in terms of fine-tuning using linear probing. It is notable that the top-performing EsViT [118] employed Swin-B [42] architecture as the backbone network. In contrast, almost all the methods obtain over 82.0% in terms of end-to-end fine-tuning. However, MSG-MAE [116], TEC [119], and dBOT [120] present the top-1 accuracy of 85.3%, 84.7%, and 84.5%, respectively.
@@
SwAV [37] ViT-B/16 300 71.6 -
SimCLR [32] ViT-B/16 300 73.9 -
BYOL [38] ViT-B/16 300 73.9 -
DINO [61] ViT-B/16 300 78.2 82.8
Data2Vec [121] ViT-B/16 800 - 84.2
MoCo-v3 [93] ViT-B/16 300 76.7 83.2
BEiT [62] ViT-B/16 800 56.7 83.2
MAE [69] ViT-B/16 800 64.4 83.4
BootMAE [122] ViT-B/16 300 64.1 84.0
MSG-MAE [116] ViT-B/16 300 75.6 85.3
TEC [119] ViT-B/16 300 - 84.7
Deit-III [123] ViT-B/16 800 - 83.8
Deit-III+AugMask [124] ViT-B/16 800 - 84.2
EsViT [118] Swin-B [42] 300 80.4 -
iBOT [117] ViT-B/16 1600 79.5 84.0
dBOT [120] ViT-B/16 800 - 84.5
CIM-RevDet [125] ViT-B/16 300 - 83.3
MFM [126] ViT-B/16 300 - 83.1
MixedAE [127] ViT-B/16 300 - 83.8
MaPeT [128] ViT-B/16 300 73.5 83.6
Table 5: Semantic segmentation ($mIoU$) comparison on ADE20K using Upernet and ViT-B networks. In addition, object detection ($AP^{box}$) and instance segmentation ($AP^{mask}$) over MS COCO datasets using Cascaded Mask R-CNN with ViT-B. The best three results are in red, green, and blue.
Methods $mIoU$ $AP^{box}$ $AP^{mask}$
@@
BEiT [62] 47.1 49.8 44.4
DINO [61] 44.1 50.1 43.4
MoCo-v3 [93] 47.3 47.9 42.7
MAE [69] 47.6 46.8 41.9
BootMAE [122] 49.1 48.5 43.4
Diet-III [123] 49.7 50.7 43.6
Diet-III+AugMask [124] 50.2 50.9 43.9
MAE [69] 48.1 50.3 44.9
MSG-MAE [116] 49.7 52.3 48.8
data2vec [121] 48.2 - -
TEC [119] 49.9 54.6 47.2
iBOT [117] 48.4 51.2 44.2
dBOT [120] 49.5 52.7 45.7
MixedAE [127] 48.9 51.0 44.1
CIM-RevDet [125] 43.5 - -
MFM [126] 48.6 - -

### v6

@@ 3.1.4 Hybrid Methods
Hybrid methods similar to generative approaches often combine the strengths of different techniques to achieve specific objectives (Shao et al. 2024).
@@
(Zhao et al. 2025), proposed a hybrid deep learning model MAEMC-NET, which combines masked autoEncoders for reconstructing masked CT patches and Momentum Contrast (MoCo) for learning discriminative features from different CT views.
(Wu and Xu 2024) implemented the hybrid SSL pre-training framework Enhanced SSL with Masked Autoencoders (ESLA). This innovative approach integrates both contrastive learning (CL) and MAE frameworks to enhance SSL. ESLA improves model generalization and representation while addressing issues stemming from competing features during data enhancement. This is achieved through a masking mechanism in the CL framework and by blending representations of different views in the latent space.
@@
Quan et al. (2022) introduced a novel SSL model, based on the global contrast-masked autoencoder (GCMAE), which adeptly extracts features from pathological images including both local and global features. GCMAE leverages masking image reconstruction and contrastive learning as self-supervised pretext tasks, enabling the encoder part to proficiently depict local and global features. This model is comprises of an initializer, encoder, a tile feature extractor, and a global feature extractor facilitating image reconstruction and contrastive learning tasks. In this investigation, an asymmetric encoder-decoder architecture is employed. ViT serves as the backbone for the encoder, pretrained using MAE. The module responsible for extracting features from patches consists of 8 transformer blocks, creating an asymmetric configuration when combined with an encoder comprising approximately 12 transformer blocks, as seen in the ViT-base architecture. This asymmetry facilitates decoupling of the encoder and decoder, fostering the acquisition of more generalized representations. Extraction of global feature is achieved via utilizing contrastive learning. The hidden representation Vpatch(vis) is used to reconstruct image through decoder and update a memory bank (B). This updation of B is executed to store the global features utilizing momentum coefficient t. The dynamically update is conducted to pursue the feature vectors of image data using memory bank as a fixed size queue. The features randomly selected from memory bank are utilized as negative samples, that helps to reduce the effects of the constraint caused by batch size on the heavy performance Notably, the momentum adjustment mechanism is designed distinctively from MoCo’s model parameter update, it employ an encoder structure and a memory bank to facilitate contrastive learning. Vpath(vis) is depicted as the encoders output in the current epoch in the latent representation, where as the vpatch(vis) represented the previous epoch latent representation stored in B. Particularly, the latent demonstrations at the output of encoder as vpatch(vis) in the current epoch, The cost function combines both, the weighted sum of MSE loss for tile feature extraction and the NCE loss for global feature extraction. This formulation aims to minimize the gap between similar features, enhancing the model’s generalization and accuracy, particularly in cross-dataset transfer learning tasks.
@@
Fang et al. (2023) proposed a hybrid SSL framework (HSL) for Hyperspectral Image (HSI) classification. Leveraging MIM and contrastive learning to enhance its effectiveness. Employing a ViT as the backbone network within an asymmetric encoder-decoder architecture with two branches, HSL achieve improved HSI classification utilizing contrastive pre-training and self-supervised generative tasks by effectively extracting spatial and spectral details. To facilitate image reconstruction through masked inputs, each branch connects intermediate features of the encoder to the decoder at the corresponding level. Contrastive learning is accomplished by utilizing the intermediate features that are obtained from two encoders that share parameters across both branches. Utilizing standard data augmentation methods, HSL generates several views of input images. Then each image view undergoes random masking technique such as MAE. Initially, image segments with mask are reorganised into a series of patches and later processed by HSL encoder and decoder to extract patterns. Extracted features are then pass through an adaptive average pooling layer and a projection head to achieve the contrastive learning task of capturing image resemblance. The decoder, which also includes multiple Transformer encoders, rebuilds patches after certain patches have been masked. Moreover, the Transformer Encoder retains the initial structure while the two output features from the encoder are used to compute contrastive loss. The loss is calculated using the decoded sequence produced by the decoder and the sequence that precedes the encoder masking, efficiently completing the masked image rebulding process. In summary, HSL combines contrastive pre-training and self-supervised generative tasks, facilitating effectiveness for downstream HSI classification task.
@@ 4.1 BYOL
Bootstrap Your Own Latent (BYOL) (Grill et al. 2020) employs a teacher-student framework, where the student network predicts the output of a teacher network on augmented inputs. The teacher network is updated as a moving average of the student’s parameters, ensuring learning consistency. This method mirrors the stabilization effect seen in regularization techniques like batch normalization and parameter sharing.
@@ 4.2 MoCo
Momentum Contrast (MoCo) (He et al. 2020) leverages a momentum-based moving average of the query encoder, establishing a large and consistent dictionary. This approach acts as a dynamic regularization mechanism, ensuring stability in the representation of dictionary keys over time. This stability is vital for the extraction of stable and generalizable features, akin to how traditional regularization methods prevent overfitting by maintaining consistency in model parameters.
@@ 4.3 SimCLR
SimCLR (Chen et al. 2020) utilizes a straightforward contrastive learning framework, where its regularization is primarily through aggressive data augmentation techniques. These augmentations consist of random cropping, resizing, color manipulation, and the application of Gaussian blur, enabling the model to learn invariant and robust features. This is similar to how techniques like dropout and data augmentation in traditional settings prevent over-reliance on specific input features.
@@ 4.4 DINO
SINO (Self-Distillation with No Labels) (Caron et al. 2021) utilizes self-distillation for regularization, where the student network mimics the teacher network, updated as an exponential moving average of the student’s parameters. This method is reminiscent of techniques like early stopping and weight decay, which aim to prevent overfitting by smoothing the learning trajectory.
@@ 4.5 Barlow Twins
Barlow Twins (Zbontar et al. 2021) method aims to align the cross-correlation matrix between the outputs of twin networks with the identity matrix. This alignment encourages the model to learn features invariant to input distortions, functioning as a regularization technique by discouraging the learning of redundant features.
@@ 4.6 SwAV
SwAV (Swapping Assignments between Views) (Caron et al. 2020) regularization is unique, involving clustering approaches and consistency enforcement between cluster assignments for alternate projections of the single image. This approach is akin to enforcing a constraint (regularization) on the feature space to enhance generalizability.
@@ 4.7 TiCo
TiCo (Time-Contrastive Consistent Learning of Visual Representations) (Zhu et al. 2022) focuses on learning temporally consistent representations, acting as a regularization mechanism by ensuring robustness to temporal variations in data. This is similar to entropy regularization, which encourages the model to learn more uncertain, hence generalizable, features.
@@ 4.8 CMC
CMC (Contrastive Multiview Coding) (Tian et al. 2020) regularizes learning by enforcing invariance to different views of the same scene. This invariance is a form of regularization, similar to dropout variations that encourage the model to learn features that are invariant to specific changes in the input.
@@ 4.9 SELF-LABELLING
SELF-LABELLING (Asano et al. 2019) approach involves pseudo-labeling, where the network assigns labels to its inputs and learns from these self-generated labels. This process regularizes the model by forcing it to generalize from its initial, potentially noisy, self-generated labels, akin to label smoothing which also deals with label noise.
@@ 5.2.1 ImageNet
The ImageNet classification dataset (Russakovsky et al. 2015) contains roughly 1.2 million images for training, fifty thousand images for validation, and 100000 images for testing. Overall, the dataset is divided into one thousand diverse object classes including tablelamp, goldfish, radiator, screwdriver, and so on.
@@ 5.2.2 COCO
COCO (Common Objects in Context) (Lin et al. 2014) is a large-scale object detection, key-point detection, segmentation, and captioning dataset. The dataset contains over 330,000 images of which around 200,000 are fully labeled. The dataset is divided into 80 object classes and over 90 generic ‘stuff’ classes.
@@ 5.2.3 CIFAR-10
The CIFAR-10 a short for Canadian Institute For Advanced Research is a dataset (Krizhevsky et al. 2009) that contains sixty thousand images divided into 10 classes in which every class contains 6000 images. In total, 50000 images are utilized for training purpose while the rest of 10000 images are utilized for testing purpose. The dataset comprises of 10 classes, including horse, airplane, frog, automobile, cat, deer, bird, dog, ship, and truck.
@@ 5.2.4 CIFAR-100
CIFAR-100 dataset (Krizhevsky et al. 2009) is an extension of CIFAR-10. Although the total number of images are same on both datasetss, i.e., sixty thousand, CIFAR-100 is divided into 100 distinct classes with 600 images per class. Similar to CIFAR-10, a total of 50000 images are utilized for training and 10000 images are employed for testing. Every image contains two labels, a coarse label (general category) and a fine label (particular class). Some particular object classes include apple, bee, cattle, chair, rabbit, and wowan, whereas general categories include flowers, household_furniture, medium_mammals, people, etc.
@@ 5.3.1 InfoNCE Loss
InfoNCE loss (Information Noise Contrastive Estimation) (Oord et al. 2018) is a loss function commonly used in contrastive learning, particularly in self-supervised learning tasks. It operates by introducing noise into the input data to create negative pairs, enabling the model to differentiate between real positive pairs and artificially generated negatives.
@@ 5.3.2 NT-Xent/SimCLR Loss
NT-Xent (Normalized Temperature Scaled Cross Entropy) (Chen et al. 2020), which extends InfoNCE by incorporating a temperature parameter to scale the pairwise similarities. Through a softmax function, it computes the negative log-likelihood of positive pairs, enhancing the overall contrastive learning process.
@@ 5.3.3 Triplet Loss
In many machine learning tasks, particularly in face recognition, person identification, and similarity-based retrieval, a key challenge is learning meaningful feature representations that ensure similar instances are clustered together while dissimilar ones are well separated. Traditional classification-based losses struggle in scenarios with high intraclass variability and low interclass differences, leading to misclassifications. Triplet loss (Schroff et al. 2015) addresses this issue by using a relative distance-based approach, which ensures that an anchor sample is closer to a positive sample (the same class) than to a negative sample (different class) by a specified margin. By simultaneously minimizing the anchor-positive distance and maximizing the anchor-negative distance, Triplet Loss enforces a structured embedding space that enhances discrimination. This approach has proven highly effective in deep metric learning tasks, such as FaceNet, where learning compact and well-separated embeddings is crucial for robust face verification, clustering, and retrieval in unseen data scenarios.
@@ 5.3.6 Circle Loss
In response to the shortcomings of triplet loss, the Circle Loss(Sun et al. 2020) was recently developed. By establishing a decision boundary in a hypersphere around instances and incorporating a flexible buffer based on the instance distances, it encourages positive pairs to lie within the boundary while keeping negative pairs outside.
These are just a few instances of the wide array of contrastive loss functions employed in various research works. The selection of a loss function is determined by factors like the specific problem, the nature of the data, and the required characteristics of the learned representations. Researchers often experiment with multiple loss functions to determine the most suitable one for their specific task (Wang et al. 2020). As contrastive learning progresses, more innovative loss formulations and improvements are likely to emerge, opening up new possibilities for unsupervised representation learning .
@@ 5.4 Comparison with Supervised Learning Approaches
SSL, particularly in the context of ViTs, has completely changed the area of computer vision. Large volumes of unlabeled data can be used with these techniques to acquire meaningful representations, reducing the substantial reliance on labeled instances observed in conventional supervised learning. The main benefit of SSL is its capacity to take advantage of the wealth of unlabeled data that is readily available online (Chen et al. 2020), allowing models to learn from big datasets without the need for explicit annotations. SSL is especially appealing for real-world applications because of its scalability and affordability given that manual data labeling can be resource-intensive, costly, and require significant effort (Gui et al. 2024).
The comparative evaluation of SSL and supervised learning methods using the same downstream tasks and evaluation metrics has shed light on the effectiveness of SSL in ViTs. In situations where the annotated data is limited or challenging to gather, SSL has demonstrated its potential to achieve competitive or even superior performance. This is especially valuable for real-world scenarios where manually annotating data might not be feasible, yet high-performance computer vision models are needed. A comprehensive learning framework known as ”SSL” employs pretext tasks derived solely from unsupervised data. These pretext tasks are designed to learn the meaningful visual representation in order to complete them effectively (Chen et al. 2020). Whereas, in the supervised learning paradigm, each input sample is connected to its matching target or output label, and the system is trained on a labeled dataset. In order for the model to accurately predict outcomes on data that has not yet been observed, supervised learning aims to establish a relationship between input data and their associated output labels (Pan and Yang 2009).
@@ 6 Comparative Analysis of SSL Mechanisms
In this section, we will provide a comparative analysis of state-of-the-art SSL (SSL) mechanisms employing ViTs. Various SSL methods used different networks as their backbone networks. For an unbiased comparison, we compare different SSL methods employing ViT-B/16 (Dosovitskiy et al. 2020) as their backbone network.
@@ 6.1 Performance Comparison
We compare SOTA SSL methods, in Table 4, where self-supervised pre-training is performed over ImageNet-1K (Russakovsky et al. 2015) training dataset. Later, fully supervised training is performed to evaluate the feature representations using (i) linear probing and (ii) end-to-end fine-tuning. The inputs are resized to 224x224 crops and report top-1 validation accuracy. From Table 4, we notice that DINO (Grill et al. 2020), MoCo-V3 (Chen* et al. 2021), MSG-MAE (Tukra et al. 2023), iBOT (Zhou et al. 2021), and EsViT (Li et al. 2021) achieve more than 75% top-1 accuracy, in terms of fine-tuning using linear probing. It is notable that the top-performing EsViT (Li et al. 2021) employed Swin-B (Liu et al. 2021) architecture as the backbone network. In contrast, almost all the methods obtain over 82.0% in terms of end-to-end fine-tuning. However, MSG-MAE (Tukra et al. 2023), TEC (Gao et al. 2022), and dBOT (Liu et al. 2022) present the top-1 accuracy of 85.3%, 84.7%, and 84.5%, respectively.
@@
SwAV (Caron et al. 2020) ViT-B/16 300 71.6 -
SimCLR (Chen et al. 2020) ViT-B/16 300 73.9 -
BYOL (Grill et al. 2020) ViT-B/16 300 73.9 -
DINO (Caron et al. 2021) ViT-B/16 300 78.2 82.8
Data2Vec (Baevski et al. 2022) ViT-B/16 800 - 84.2
MoCo-v3 (Chen* et al. 2021) ViT-B/16 300 76.7 83.2
BEiT (Bao et al. ) ViT-B/16 800 56.7 83.2
MAE (He et al. 2022) ViT-B/16 800 64.4 83.4
BootMAE (Dong et al. 2022) ViT-B/16 300 64.1 84.0
MSG-MAE (Tukra et al. 2023) ViT-B/16 300 75.6 85.3
TEC (Gao et al. 2022) ViT-B/16 300 - 84.7
Deit-III (Touvron et al. 2022) ViT-B/16 800 - 83.8
Deit-III+AugMask (Heo et al. 2023) ViT-B/16 800 - 84.2
EsViT (Li et al. 2021) Swin-B (Liu et al. 2021) 300 80.4 -
iBOT (Zhou et al. 2021) ViT-B/16 1600 79.5 84.0
dBOT (Liu et al. 2022) ViT-B/16 800 - 84.5
CIM-RevDet (Fang et al. 2022) ViT-B/16 300 - 83.3
MFM (Xie et al. 2022) ViT-B/16 300 - 83.1
MixedAE (Chen et al. 2023) ViT-B/16 300 - 83.8
MaPeT (Baraldi et al. 2023) ViT-B/16 300 73.5 83.6
Table 5: Semantic segmentation (mIoUmIoU) comparison on ADE20K using Upernet and ViT-B networks. In addition, object detection ($AP^{box}$) and instance segmentation ($AP^{mask}$) over MS COCO datasets using Cascaded Mask R-CNN with ViT-B. The best three results are in red, green, and blue.
Methods mIoUmIoU $AP^{box}$ $AP^{mask}$
@@
BEiT (Bao et al. ) 47.1 49.8 44.4
DINO (Caron et al. 2021) 44.1 50.1 43.4
MoCo-v3 (Chen* et al. 2021) 47.3 47.9 42.7
MAE (He et al. 2022) 47.6 46.8 41.9
BootMAE (Dong et al. 2022) 49.1 48.5 43.4
Diet-III (Touvron et al. 2022) 49.7 50.7 43.6
Diet-III+AugMask (Heo et al. 2023) 50.2 50.9 43.9
MAE (He et al. 2022) 48.1 50.3 44.9
MSG-MAE (Tukra et al. 2023) 49.7 52.3 48.8
data2vec (Baevski et al. 2022) 48.2 - -
TEC (Gao et al. 2022) 49.9 54.6 47.2
iBOT (Zhou et al. 2021) 48.4 51.2 44.2
dBOT (Liu et al. 2022) 49.5 52.7 45.7
MixedAE (Chen et al. 2023) 48.9 51.0 44.1
CIM-RevDet (Fang et al. 2022) 43.5 - -
MFM (Xie et al. 2022) 48.6 - -